Can yo help me to do this problem please! : It takes a plane 5 hours to travel 3000 miles with a tailwind, but it takes tha same plane 6 hours to travel the same distance with a headwind. What is the speed of the plane in still air and the speed of tha wind? Write an equation.
Answer by checkley71(8403)

DISTANCE=RATE TIMES TIME
3000=R*5
R=3000/5
R=600 MPH WITH A TAIL WIND
3000=R*6
R=3000/6
R=500 MPH
PROOF
600-W=500+W
100=2W
W=100/2
W=50 MPH THE SPEED OF THE WIND